OpenStudy (anonymous):

solve (3x+7)^1/2=x-1

OpenStudy (anonymous):

6

OpenStudy (gw2011):

(3x+7)^1/2 = x-1 Square both sides since (3x+7)^1/2 is a square root 3x+7 = (x-1)^2 3x+7 = x^2-2x+1 0 = x^2-5x-6 I will rewrite this equation to get: x^2-5x-6 = 0 (x-6)(x+1) = 0 x-6 = 0 x = 6 x+1 = 0 x = -1 The two possible solutions are x=6 and x=-1 Now you need to substitute these values to see what the solution really is [(3)(6)+7]^1/2 = 6-1 (18+7)^1/2 = 6-1 (25)^1/2 = 6-1 5 = 5 which is a true statement and, therefore, x=6 is true (valid) [(3)(-1)+7]^1/2 = -1-1 (-3+7)^1/2 = -1-1 (4)^1/2 = -1-1 2 is not equal to -2 and, therefore, x=-1 is not true (not valid) Therefore, the answer is x=6